OpenStudy (anonymous):

What are the roots of 2x+6=−5x? 3−+i2 −3−+i2 −3−+2i4 3−+i2‾‾√4 please help

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@t.ashleyxx

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@dom4958

OpenStudy (mathstudent55):

Shouldn't one of the x's be squared?

OpenStudy (anonymous):

ill attach a screenshot it doesn't post right

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

That's a completely different problem.

OpenStudy (anonymous):

can you help?

OpenStudy (mathstudent55):

With the screenshot?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

We need to multiply this together. \((1 + i\sqrt 3)(1 + 2i\sqrt 3) \) We use FOIL. \(=1 + 2i\sqrt 3 + i \sqrt 3 + 2i^2 (\sqrt 3)^2\) Ok so far?

OpenStudy (mathstudent55):

\(=1 + 3i\sqrt 3 - 6\) \(=-5 + 3i\sqrt 3 \)

OpenStudy (anonymous):

okay